log (5x+1) base 4 = 2
Answer by stanbon(75887)

log (5x+1) base 4 = 2
----
log4(5x+1) = 2
----
base = 4
log = exponent = 2
result = 5x+1
-----
Solve:
5x+1 = 4^2
5x+1 = 16
5x = 15
x = 3
---------------
Cheers,
Stan H.
==========